Feed aggregator

HYST AREA 6-2-1 Cycles Magnets

Recent Log Entries -

Log Number: 3818978Logbook(s): Body: 
HYST AREA Version 6-2-1 has FAILED to cycle:

MBDAS10H -- W102B10 -- wa1mag -- Failed to set after 2 cycle attempts


HYST AREA Version 6-2-1 has successfully cycled the following:

Area.Arcs.Arc2.Spreader
Area.Arcs.Arc2.Arc
Area.Arcs.Arc2.Recombiner
Area.Arcs.Arc2.Extraction
Area.Arcs.Arc4.Spreader
Area.Arcs.Arc4.Arc
Area.Arcs.Arc4.Recombiner
Area.Arcs.Arc4.Extraction
Area.Arcs.Arc6.Spreader
Area.Arcs.Arc6.Arc
Area.Arcs.Arc6.Recombiner
Area.Arcs.Arc6.Extraction
Area.Arcs.Arc8.Spreader

Box Supplies Were Cycled

Recent Log Entries -

Log Number: 3818977Logbook(s): Tags: Body: 

The following box supplies were CYCLED by version 2-7 of the Box Supply Cycling Script:
ARC2A Ramp Rate = 6.895 A/s
ARC4A Ramp Rate = 11.048 A/s
ARC6A Ramp Rate = 9.130 A/s
ARC8A Ramp Rate = 11.948 A/s
ARCAA Ramp Rate = 12.239 A/s
RSEP8A Ramp Rate = 10.471 A/s

Day shift summary 8/1/2020

Recent Log Entries -

Log Number: 3818970Logbook(s): Tags: Entry Makers: aruntsBody: 

Shift summary 8/1/2020

1. Issue with HMS dipole NMR lock - Steve Lassiter said he will take a look at it after lunch.
2. Mingyu finished target tests (AFP loss studies, PC temperature study, NMR cross calibration with different target ladder positions and pNMR vs NMR calibration)
3. SHMS Q1, Q2, Q3, HB and dipole set.
4. Simona looked at cosmics and opened up the timing cuts.

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Recent Log Entries -

Log Number: 3818958Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: gauthier
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09
ILM4C01
ILM4C02
ILM4C03
ILM1C

Downtime Incident CLOSED: 1L14-4 SOSing

Recent Log Entries -

Log Number: 3818949Logbook(s): Body: IDDescriptionPeriodCause8266Title:
1L14-4 SOSing

Summary:
GSET was 7.362. Tried retuning; noticed RATN had to be 0 to get good forward power from GLOD. Paged EESRF for help as we try avoiding Gradient loss. EESRF is working on it remotely.

1L14-4 Low Drive

Recent Log Entries -

Log Number: 3818942Logbook(s): Entry Makers: kernsBody: 

GASK at 6.7v at 4Mv/m. No more RATN, it's at 0.0 V. The cavity starts SOSing if gradient is raised and GASK rails goes to 9.9v . That means the 2.5w amp or RFCM output is low. Most of the time the Amp. Sam and I will come in to replace the amp since its the North and needed to keep trip rates down. GASK at 6.6v with GMES of 4Mv/m . It is stable but won't run any higher.

Pages

Subscribe to cebaf2013 aggregator